On january 1, 2016 Charles Jemison borrows $40000 from his father to open a business . Charles is the beneficiary of a trust created by his aunt from which he will receive $25000 on january 1 2016 . He signs an agreement to make this amount payable to his father and, further ,to pay his father equal annual amounts from january 1, 2017   to january 1, 2025 , inclusive , in retirement of the debt. Interest is 12% per year
Answer #1

Total amount pending to be paid by Charles to his father=$40000-$25000=$15000

From 2017 to 2025 ,9 instalments have to be made at $1667 per instalments

Calculation of interest on outstanding balance of debt

  

Date Instalments Outstanding Interest payment
Jan 1 2017 1667 15000 15000*12/100=1800
Jan 1 2018 1667 15000-1667=13333 1600
Jan 1 2019 1667 13333-1667=11666 1400
Jan 1 2020 1667 11666-1667=9999 1200
Jan 1 2021 1667 9999-1667=8332 1000
Jan 1 2022 1667 8332-1667=6665 800
Jan 1 2023 1667 6665-1667=4998 600
Jan 1 2024 1667 4998-1667=3331 400
Jan 1 2025 1664 3331-1667=1664 200


Total payment

Years Total payment
2017 1667+1800=3467
2018 1667+1600=3267
2019 3067
2020 2867
2021 2667
2022 2467
2023 2267
2024 2067